LINE INTENSITY MEASUREMENTS IN ^{14}N_{2}^{16}O AND THEIR TREATMENT USING THE EFFECTIVE OPERATOR APPROACH. II. THE 5200 TO 6400 CM^{−1} REGION

\underline{J. VANDER AUWERA}, Laboratoire de Chimie Physique Moléculaire C. P. 160/09, Université Libre de Bruxelles, 50 Avenue F. D. Roosevelt, B-1050 Brussels, Belgium; L. DAUMONT, Groupe de Spectrométrie Moléculaire et Atmosphérique, ESA CNRS 6089, Université de Reims, Faculté des Sciences, Moulin de la Housse, BP 1039, F-51687 Reims cedex 2, France; V. I. PEREVALOV, S. A. TASHKUN, Institute of Atmospheric Optics, Russian Academy of Sciences, Siberian Branch, 1 Akademicheskii Avenue, 634055 Tomsk, Russia; J-L. TEFFO, Laboratoire de Physique Moléculaire et Applications, CNRS, boîte 76, Université Pierre et Marie Curie, 4 Place Jussieu, F-75252 Paris, France.

This work continues a series of publications devoted to the application of the effective operators approach to the analysis and prediction of vibration-rotation spectra of linear triatomic molecules. In that frame, the present work aims at describing line intensities of cold and hot bands of ^{14}N_{2}^{16}O in its ground electronic state in the spectral range above 3600 cm^{−1}. In N_{2}O, vibrational interacting levels group in polyads, identified by the so-called polyad number P = 2V_{1}+V_{2}+4V_{3}, as a result of the relation 2ω_{1} ∼ 4ω_{2} ∼ ω_{3} existing between the harmonic frequencies. The absorption spectra of N_{2}O, at room temperature, have been recorded in Brussels over the whole range between 3600 and 11000 cm^{−1} using a Bruker IFS120HR Fourier transform spectrometer. The measurement and analysis of absolute line intensities in the region between 4300 and 5200 cm^{−1}, involving bands associated with transitions corresponding to ΔP = 7, 8 and 9, was done recently.^{a} We are now measuring absolute line intensities for cold and hot bands associated with transitions corresponding to ΔP = 10 and 11, observed in the range from 5200 to 6400 cm^{−1}. Using wavefunctions previously determined from a global fit of an effective hamiltonian to about 18000 line positions,^{b} parameters of a corresponding effective dipole moment are then fitted to these experimental intensities. Results will be presented and discussed.

aL. Daumont, J. Vander Auwera, J-L. Teffo, V. I. Perevalov and S. A. Tashkun, J. Mol. Spectrosc. \underline{208}, 281-291 (2001).

bA. Camparque, G. Weirauch, S. A. Tashkun, V. I. Perevalov, and J-L. Teffo, J. Mol. Spectrosc. \underline{209}, 198-206 (2001).